# Completely regular growth solutions to linear differential equations with exponential polynomials coefficients

### v1

@@ 3.1 LDE with exponential sums coefficients
$\sum_{j=0}^{N_{0}}e^{\lambda_{j}z}\left(a_{n,j}(z)f^{(n)}(z)+a_{n-1,j}(z)f^{(n-1)}(z)+\cdots+a_{0,j}(z)f(z)\right)=0,\quad$ (3.3)
@@
are called coefficient differential equations about $\lambda_{j}$ (or at $\overline{\lambda}_{j}$) individually for $j=0,\ldots,N_{0}$. Set $W=\left\{\overline{\lambda}_{j}\right\}$, the convex hull $co(W)$ of $W$ is defined as the intersection of all closed convex sets containing $W$, also to be said as the convex hull to the differential equation (3.2). Suppose the vertexes of $co(W)$ consist a set $\tilde{W}=\{\overline{\omega}_{j}\}(j=0,\ldots,N_{0}^{\prime})$ with elements arranged anticlockwise, $N_{0}^{\prime}\leq N_{0}$, and denote $s_{j}$ the segment between $\overline{\omega}_{j}$ and $\overline{\omega}_{j+1}$. The critical ray $\arg z=\eta_{j}$ is defined as the ray originated at 0 with the direction of the outer normal to $s_{j}$. For small enough $\varepsilon>0,$ we define
$S_{j}(\varepsilon)=\{z|\eta_{j-1}+\varepsilon\leq\arg z\leq\eta_{j}-\varepsilon\},\enspace T_{j}(\varepsilon)=\{z|\eta_{j}-\varepsilon<\arg z<\eta_{j}+\varepsilon\}$
@@
    $0\leq\arg z-\xi_{jk}<\frac{\log^{+}|z|}{{|z|}^{\frac{1}{p_{j}}}},k,p\in\mathbb{N^{*}}$
@@
Note. The polynomial $G_{j,\theta}$ will change according to the rotation of $\arg z=\theta$ from one side of $S_{j}$ to the other inside $S_{j}$, but this change occurs only when $\arg z=\theta$ is a Stokes ray in $S_{j},j=1,\ldots,N_{0}^{\prime}$.
@@
$G_{1}(g)=\frac{g^{\prime\prime\prime}}{g}-\frac{4}{3}\frac{g^{\prime}}{g}+\frac{16}{27},\enspace\text{G}_{2}(g)=3\frac{g^{\prime\prime}}{g}-2\frac{g^{\prime}}{g}-1,$
@@
$G_{1}(f(z))=1-\frac{1}{1-7e^{z}}\rightarrow 0,\enspace z\in S_{0}(\varepsilon)=\{z|\pi/2+\varepsilon\leq\arg z\leq 3\pi/2-\varepsilon\},$
@@
$G_{2}(f(z))=\frac{7}{1-7e^{z}}\rightarrow 0,\enspace z\in S_{1}(\varepsilon)=\{z|-\pi/2+\varepsilon\leq\arg z\leq\pi/2-\varepsilon\}.$
@@
$04ze^{iz}f^{\prime\prime\prime}+(6e^{iz}+4ize^{iz}-z)f^{\prime\prime}+\left((2i-1)e^{iz}-\frac{1}{2}\right)f^{\prime}-\left(ie^{iz}-\frac{1}{4}\right)f=$ 0 (3.4)
@@
$04zf^{\prime\prime\prime}+(6+4iz)f^{\prime\prime}+(2i-1)f^{\prime}-if=0,\quad-zf^{\prime\prime}-\frac{1}{2}f^{\prime}+\frac{1}{4}f=$ 0 .
@@ 3.2 LDE with exponential polynomials coefficients
$G_{m}(z)=\sum_{t=1}^{l_{m}}P_{m,t}(z)\exp(Q_{m,t}(z)),\enspace 0\leq m\leq n-1,\enspace l_{m}\in\mathbb{N},$ (3.7)
@@
where $k$ is the maximal number of pairwise different complex number $q_{j},$ appearing in all coefficients of $z^{s}$ in $Q_{m,t}$, $q_{1}=0$ and $s=\max\{\deg(Q_{m,t})\}$. Thereby, the degree of every polynomial $\hat{G}_{m,j}(z)$ is smaller or equal to $s-1.$
@@
$S_{j_{1}}(\varepsilon)=\{z|\eta_{j_{1}-1}+\varepsilon\leq\arg z\leq\eta_{j_{1}}-\varepsilon\},\enspace T_{j_{1}}(\varepsilon)=\{z|\eta_{j_{1}}-\varepsilon<\arg z<\eta_{j_{1}}+\varepsilon\},$
@@
In Figure 2, $n(\cdot)$ is a number controlled by the variables ``"``"`$\cdot$"` ` ⋅ ". To write inductively, we set $n(j_{1},\ldots,j_{t-1})=k_{j_{1},\ldots,j_{t-1}}$. We have already separated the dominant exponential polynomial terms $e^{q_{j_{1}}z^{s}}$ in (3.8) and their coefficient differential equation (3.9) at $q_{j_{1}}$. Within each coefficient differential equation (3.9) at $q_{j_{1}}$, we can do this transformation again and sparate the dominant exponential polynomial terms $e^{q_{j_{1},j_{2}}z^{s-1}}$ and their coefficient differential equations at $q_{j_{1},j_{2}}$, where $q_{j_{1},j_{2}}$ are leading coefficients to the degree $s-1$. We denote the convex hull to equation (3.9) as $co(W_{j_{1}}),j_{1}=1,\ldots,k$; denote the $r_{j_{1},j_{2}}$ many conjugated leading coefficients on the vertexes of convex hull $co(W_{j_{1}})$ as $\overline{\omega}_{j_{1},j_{2}},j_{1}=1,\ldots,k,j_{2}=1,\ldots,k_{j_{1}}$ in counterclockwise order, where $k_{j_{1}}$ is the maximal number of pairwise different complex number $q_{j_{1},j_{2}}$; denote the $r_{j_{1},j_{2}}$ many critical rays $\arg z=\eta_{j_{1},j_{2}},$ defined as the ray originated at 0 with the direction of the outer normal to the segment between $\overline{\omega}_{j_{1},j_{2}}$ and $\overline{\omega}_{j_{1},j_{2}+1}$, corresponding sectors as
$S_{j_{1},j_{2}}(\varepsilon)=\{z|\eta_{j_{1},j_{2}-1}+\varepsilon\leq\arg z\leq\eta_{j_{1},j_{2}}-\varepsilon\},\enspace T_{j_{1},j_{2}}(\varepsilon)=\{z|\eta_{j_{1},j_{2}}-\varepsilon<\arg z<\eta_{j_{1},j_{2}}+\varepsilon\},$
@@
$\overline{\omega}_{j_{1},\ldots,j_{t}+1}$$\overline{\omega}_{j_{1},\ldots,j_{t}}$$\overline{\omega}_{j_{1},\ldots,j_{t}-1}$$\overline{\omega}_{j_{1},\ldots,j_{t-1},1}$$\overline{\omega}_{j_{1},\ldots,j_{t-1},r_{j_{1},\ldots,j_{t-1}}}$${co(W_{j_{1},\ldots,j_{t-1}})}$$\arg z=\eta_{j_{1},\ldots,j_{t}}$$\varepsilon$$\varepsilon$$\varepsilon$$S_{j_{1},\ldots,j_{t}}(\varepsilon)$$T_{j_{1},\ldots,j_{t}}(\varepsilon)$$\arg z=\eta_{j_{1},\ldots,j_{t}-1}$$\eta_{j_{1},\ldots,j_{t-1},1}$$\eta_{j_{1},\ldots,j_{t}+1}$$\eta_{j_{1},\ldots,j_{t-1},r_{j_{1},\ldots,j_{t-1}}}$
@@
Additionally, we need some relations about the distribution of $z$ and $z^{s}$, where $s\in\mathbb{C}$. Let $10\leq\alpha<$ 1 and
@@ 4 Proofs
By Gundersen [4, Corollary 3], if we set $\rho=\rho(f)$ as growth order of $f$, there exists $\varepsilon_{0}>0$ and a set $E$ with finite linear measure, such that
$\left|\frac{f^{(k)}}{f}\right|<{|z|^{(k-1)(\rho+\varepsilon_{0})}},\quad z\notin E.$
@@
$\sum\limits_{t=0}^{n}a_{t,j}\dfrac{f^{(t)}}{f}=o\left(\frac{1}{z^{N_{1}}}\right),\enspace\text{for any}\enspace N_{1}\in\mathbb{N},\enspace z\in S_{j}(\varepsilon),\enspace|z|\notin E.$ (4.3)
@@
$f^{(n)}+P_{n-1,j}f^{(n-1)}+\cdots+P_{1,j}f^{\prime}+\left(P_{0,j}+o\left(\frac{1}{z^{N_{1}+\deg(a_{0,j})}}\right)\right)f=0,$ (4.5)
@@
$P_{i,j}(z)=z^{d_{i,j}}\sum_{v=0}^{\infty}\frac{b_{i,j,v}}{z^{v}},\quad|z|>|z_{0}|,$
for $d_{i,j}\in\mathbb{N},b_{i,j,v}\in\mathbb{C},i=1,\ldots,n-1,$ $j=0,\ldots,N_{0}^{\prime},$ enough large $|z_{0}|$, $z\in S_{j}(\varepsilon),$ and $|z|\notin E$. $Q_{0}$ is consequently in the asymptotic form of
$Q_{0,j}(z)\sim z^{d_{0,j}}\sum_{v=0}^{\infty}\frac{b_{0,j,v}}{z^{v}},\quad|z|>|z_{0}|,$
@@
It is obtained that (4.5) is in the form of (2.9). Using Lemma 2.4, there exists a polynomial $G_{\theta}$ in $z^{1/p_{j}}$ and a a set of zero relative measure $E_{\theta}$ such that

### v2

@@ 3.1 LDE with exponential sums coefficients
$\sum_{j=1}^{N_{0}}e^{\lambda_{j}z}\left(a_{n,j}(z)f^{(n)}(z)+a_{n-1,j}(z)f^{(n-1)}(z)+\cdots+a_{0,j}(z)f(z)\right)=0,$ (3.3)
@@
are called coefficient differential equations about $\lambda_{j}$ (or at $\overline{\lambda}_{j}$) individually for $j=1,\ldots,N_{0}$. Set $W=\left\{\overline{\lambda}_{j}\right\}$, the convex hull $co(W)$ of $W$ is defined as the intersection of all closed convex sets containing $W$, also to be said as the convex hull to the differential equation (3.2). Suppose the vertexes of $co(W)$ consist a set $\tilde{W}=\{\overline{\omega}_{j}\}(j=1,\ldots,N_{0}^{\prime})$ with elements arranged anticlockwise, $N_{0}^{\prime}\leq N_{0}$, and denote $s_{j}$ the segment between $\overline{\omega}_{j}$ and $\overline{\omega}_{j+1}$. The critical ray $\arg z=\eta_{j}$ is defined as the ray originated at 0 with the direction of the outer normal to $s_{j}$. For small enough $\varepsilon>0,$ we define
$S_{j}(\varepsilon)=\{z|\eta_{j-1}+\varepsilon\leq\arg z\leq\eta_{j}-\varepsilon\},\kern 5.0ptT_{j}(\varepsilon)=\{z|\eta_{j}-\varepsilon<\arg z<\eta_{j}+\varepsilon\}$
@@
    $0\leq\arg z-\xi_{jk}<\frac{\log^{+}|z|}{{|z|}^{\frac{1}{p_{j}}}},k,p\in\mathbb{N^{*}}$ (3.4)
@@
Note.
  1. 1.
    The polynomial $G_{j,\theta}$ will change according to the rotation of $\arg z=\theta$ from one side of $S_{j}$ to the other inside $S_{j}$, but this change occurs only when $\arg z=\theta$ is a Stokes ray in $S_{j},j=1,\ldots,N_{0}^{\prime}$.
  2. 2.
    Steinmetz [34] indicated us that the exceptional set in the form of logarithmic semi-strips does not affect c.r.g., since this strip does not contain any ray except the boundary ray, the Stokes ray. Namely, if there exists a $\theta_{0}>0$, and outside the strip (3.4) $f$ satisfies c.r.g. on the set of rays
    $R_{\mathbb{M}}=\{re^{i\theta}:r>0,r\notin\displaystyle\bigcup_{\theta\in\mathbb{M}\subset(\xi_{jk},\xi_{jk}+\theta_{0})}E_{\theta}\}\subset\displaystyle\bigcup_{\theta\in\mathbb{M}\subset(\xi_{jk},\xi_{jk}+\theta_{0})}R_{\theta},$
    where $R_{\theta}=\{re^{i\theta},r\notin E_{\theta},r>0\}$, $E_{\theta}$ is a $E_{0}$-set depending on $\theta$ (see the definitions in Section 2.1), then $f$ is still of c.r.g. on any ray $\theta\in\mathbb{M}\subset\left(\xi_{jk},\xi_{jk}+\theta_{0}\right)$. Then, by the compactification of $\left(\xi_{jk},\xi_{jk}+\theta_{0}\right)$, Lemma 2.1, $f$ is of c.r.g. on $R_{[\xi_{jk},\xi_{jk}+\theta_{0}]}$.
  3. 3.
    To obtain an estimate of $\log|f|$ on $T_{j}(\varepsilon)$, we need to perform a more detailed characterization of the convex hull $co(W)$ of $W=\{\overline{\lambda}_{j}\}$ to (3.3). Since $T_{j}(\varepsilon)$ is a narrow sector between $S_{j}(\varepsilon)$ and $S_{j+1}(\varepsilon)$, and each $S_{j}(\varepsilon)$ is dominated by the vertex $\overline{\omega}_{j}$, it is natural to consider whether $\log|f|$ in $T_{j}(\varepsilon)$ is dominated by both vertices $\overline{\omega}_{j}$ and $\overline{\omega}_{j+1}.$ Here we brief introduce, $\log|f|$ in $T_{j}(\varepsilon)$ might be dominated by all the elements of $W$ on the side $s_{j}$ between $\overline{\omega}_{j}$ and $\overline{\omega}_{j+1}$.
In fact, [7, Lem. 8] inspires us the degeneration rules of the coefficients $A_{m}(z)$, $m=0,\ldots,n-1$, to (3.2) in $T_{j}(\varepsilon)$. From the following conclusion, it can be seen that only the conjugate indicators which are on the boundary $s_{j}$ control the growth of these coefficients $A_{m}(z)$ in $T_{j}(\varepsilon)$.
Corollary 3.3
The normal form (3.3) of (3.2) in $T_{j}(\varepsilon)$ has an asymptotic form of
$\sum_{\overline{\lambda}_{k}\in s_{j}}e^{\lambda_{k}z}\sum_{m=0}^{n}c_{m,k}z^{d_{m,k}}(1+\varepsilon(z))f^{(m)}(z)=0,$ (3.5)
where $c_{m,k}z^{d_{m,k}}$ is the highest-order term of the polynomial $a_{m,k}(z)$ in (3.3). The side $s_{j}$ of $co(W)$ has at least two boundary points $\overline{\omega}_{j}$ and $\overline{\omega}_{j+1}.$
Consider $\overline{\lambda}_{1},\ldots,\overline{\lambda}_{K}\in s_{j}$ with $j$ fixed. Let $\tau_{k,m}=\overline{\lambda}_{k}+id_{m,k}e^{i\eta_{j}}$, where $d_{m,k}$ is the degree of $a_{m,k}(z)$, $\arg z=\eta_{j}$ is the critical ray inside $T_{j}(\varepsilon)$. Let $\Omega_{j}$ be the smallest closed convex set containing $\overline{\omega}_{j},\overline{\omega}_{j+1}$ and all $\tau_{k,m}$. Then, without generality, we assume the vertices of $\Omega_{j}$ is anticloskwise set by $\tilde{\omega}_{1},\ldots,\tilde{\omega}_{K^{\prime}}$, $\overline{\omega}_{j+1},$ and $\overline{\omega}_{j}$, where $\tilde{\omega}_{k}=\overline{\lambda}_{k}+i\max_{m}\{d_{m,k}\}e^{i\eta_{j}}$, $\overline{\omega}_{j}=\overline{\lambda}_{1}$, $\overline{\omega}_{j+1}=\overline{\lambda}_{K^{\prime}}$, and $2\leq K^{\prime}\leq K.$ Set the slope of the segment $\tilde{s}_{k}=[\tilde{\omega}_{k},\tilde{\omega}_{k+1}]$ as
$\mu_{k}=\frac{\max_{m}\{d_{m,k}\}-\max_{m}\{d_{m,k+1}\}}{|\overline{\lambda}_{k}-\overline{\lambda}_{k+1}|},$
and two strips
$\begin{split}U_{1}=&\{z|\Im(ze^{-i\eta_{j}})\geq 0,\Re(ze^{-i\eta_{j}})+\mu_{1}\log|z|\geq\delta\},\\ U_{K^{\prime}}=&\{z|\Im(ze^{-i\eta_{j}})\geq 0,\Re(ze^{-i\eta_{j}})+\mu_{K^{\prime}}\log|z|\leq-\delta\},\end{split}$
for $\delta>0$. Since when $z\to\infty$, $\mathbb{C}\backslash(U_{1}\cup U_{K^{\prime}})$ cannot cover any sector or ray (except for $\arg z=\eta_{j}$), it is sufficient for us to consider the growth of functions in $T_{j}(\varepsilon)\cap U_{1}$ and $T_{j}(\varepsilon)\cap U_{K^{\prime}}$ to get c.r.g. in $T_{j}(\varepsilon)$, outside an exceptional set that does not contain any sector. For example, we consider a trivial case – if $K^{\prime}=2$ and $\mu_{1}=0$ (More specifically, all $d_{m,k}$ are zero for $\overline{\lambda}_{k}\in s_{j}$), then $T_{j}(\varepsilon)\backslash(U_{1}\cup U_{2})$ is a strip parallel to the critical ray $\arg z=\eta_{j}$ and having it as its axis of symmetry, with a width of $2H$. A typical conclusion is given by [6, Lemma 10], applied to the result below.
Corollary 3.4
Given the symbols as defined above. Let $z\in U_{1}.$ If $\tau_{p,m}\notin\tilde{s}_{1}$, there exists some $v>0$, then
$|z^{d_{m,p}}e^{\lambda_{p}z}|<|z|^{-v}\cdot\left|z^{\max_{m}\{d_{m,1}\}}e^{\omega_{j}z}\right|.$
If $\tau_{p,m}\in\tilde{s}_{1}$, then
$|z^{d_{m,p}}e^{\lambda_{p}z}|\leq|z^{\max_{m}\{d_{m,1}\}}e^{\omega_{j}z}|\cdot\exp{(-\delta|\lambda_{p}-\omega_{j}|)}.$
It is similar to the case $z\in U_{K^{\prime}}$ by substituting $d_{m,1}$, $\omega_{j}$, and $\tilde{s}_{1}$ into $d_{m,K^{\prime}}$, $\omega_{j+1}$, and $\tilde{s}_{K^{\prime}-1}$ individually.
By Corollary 3.3 and 3.4, we can deal with the case $z\in T_{j}(\varepsilon)$ in Theorem 3.2.
Theorem 3.5
All finite order transcendental entire solutions of (3.2) are of c.r.g..
@@
$G_{1}(g)=\frac{g^{\prime\prime\prime}}{g}-\frac{4}{3}\frac{g^{\prime}}{g}+\frac{16}{27},\kern 5.0pt\text{G}_{2}(g)=3\frac{g^{\prime\prime}}{g}-2\frac{g^{\prime}}{g}-1,$
@@
$G_{1}(f(z))=1-\frac{1}{1-7e^{z}}\rightarrow 0,\kern 5.0ptz\in S_{0}(\varepsilon)=\{z|\pi/2+\varepsilon\leq\arg z\leq 3\pi/2-\varepsilon\},$
@@
$G_{2}(f(z))=\frac{7}{1-7e^{z}}\rightarrow 0,\kern 5.0ptz\in S_{1}(\varepsilon)=\{z|-\pi/2+\varepsilon\leq\arg z\leq\pi/2-\varepsilon\}.$
@@
$4ze^{iz}f^{\prime\prime\prime}+(6e^{iz}+4ize^{iz}-z)f^{\prime\prime}+\left((2i-1)e^{iz}-\frac{1}{2}\right)f^{\prime}-\left(ie^{iz}-\frac{1}{4}\right)f=0$ (3.6)
@@
$4zf^{\prime\prime\prime}+(6+4iz)f^{\prime\prime}+(2i-1)f^{\prime}-if=0,\quad-zf^{\prime\prime}-\frac{1}{2}f^{\prime}+\frac{1}{4}f=0.$
@@ 3.2 LDE with exponential polynomials coefficients
$G_{m}(z)=\sum_{t=1}^{l_{m}}P_{m,t}(z)\exp(Q_{m,t}(z)),\kern 5.0pt0\leq m\leq n-1,\kern 5.0ptl_{m}\in\mathbb{N},$ (3.9)
@@
where $k$ is the maximal number of pairwise different complex number $q_{j},$ appearing in all coefficients of $z^{s}$ in $Q_{m,t}$, $q_{1}=0$ and $s=\max\{\deg(Q_{m,t})\}$. Thereby, the degree of every polynomial $\hat{G}_{m,j}(z)$ is smaller or equal to s1.s-1.
@@
$S_{j_{1}}(\varepsilon)=\{z|\eta_{j_{1}-1}+\varepsilon\leq\arg z\leq\eta_{j_{1}}-\varepsilon\},\kern 5.0ptT_{j_{1}}(\varepsilon)=\{z|\eta_{j_{1}}-\varepsilon<\arg z<\eta_{j_{1}}+\varepsilon\},$
@@
In Figure 2, $n(\cdot)$ is a number controlled by the variables $\cdot$". To write inductively, we set $n(j_{1},\ldots,j_{t-1})=k_{j_{1},\ldots,j_{t-1}}$. We have already separated the dominant exponential polynomial terms $e^{q_{j_{1}}z^{s}}$ in (3.10) and their coefficient differential equation (3.11) at $q_{j_{1}}$. Within each coefficient differential equation (3.11) at $q_{j_{1}}$, we can do this transformation again and sparate the dominant exponential polynomial terms $e^{q_{j_{1},j_{2}}z^{s-1}}$ and their coefficient differential equations at $q_{j_{1},j_{2}}$, where $q_{j_{1},j_{2}}$ are leading coefficients to the degree s1s-1. We denote the convex hull to equation (3.11) as $co(W_{j_{1}}),j_{1}=1,\ldots,k$; denote the $r_{j_{1},j_{2}}$ many conjugated leading coefficients on the vertexes of convex hull $co(W_{j_{1}})$ as $\overline{\omega}_{j_{1},j_{2}},j_{1}=1,\ldots,k,j_{2}=1,\ldots,k_{j_{1}}$ in counterclockwise order, where $k_{j_{1}}$ is the maximal number of pairwise different complex number $q_{j_{1},j_{2}}$; denote the $r_{j_{1},j_{2}}$ many critical rays $\arg z=\eta_{j_{1},j_{2}},$ defined as the ray originated at 0 with the direction of the outer normal to the segment between $\overline{\omega}_{j_{1},j_{2}}$ and $\overline{\omega}_{j_{1},j_{2}+1}$, corresponding sectors as
$S_{j_{1},j_{2}}(\varepsilon)=\{z|\eta_{j_{1},j_{2}-1}+\varepsilon\leq\arg z\leq\eta_{j_{1},j_{2}}-\varepsilon\},\kern 5.0ptT_{j_{1},j_{2}}(\varepsilon)=\{z|\eta_{j_{1},j_{2}}-\varepsilon<\arg z<\eta_{j_{1},j_{2}}+\varepsilon\},$
@@
[…]
@@
Additionally, we need some relations about the distribution of $z$ and $z^{s}$, where $s\in\mathbb{C}$. Let $0\leq\alpha<1$ and
@@ 4 Proofs of main results
By Gundersen [13, Corollary 3], if we set $\rho=\rho(f)$ as growth order of $f$, there exists $\varepsilon_{0}>0$ and a set $E$ with finite linear measure, such that
$\left|\frac{f^{(k)}}{f}\right|<{|z|^{k(\rho+\varepsilon_{0})}},\quad|z|\notin E.$
We denote
$L_{j}(f)=\sum\limits_{t=0}^{n}a_{t,j}\dfrac{f^{(t)}}{f},$
so (4.1) is transformed to
$\sum_{j=0}^{N_{0}}e^{\lambda_{j}z}L_{j}(f)=0.$
@@
$L_{j}(f)=o\left(\frac{1}{z^{N_{1}}}\right),\kern 5.0pt\text{for any}\kern 5.0ptN_{1}\in\mathbb{N},\kern 5.0ptz\in S_{j}(\varepsilon),\kern 5.0pt|z|\notin E.$ (4.3)
@@
$f^{(n)}+P_{n-1,j}f^{(n-1)}+\cdots+P_{1,j}f^{\prime}+\left(P_{0,j}+o\left(\frac{1}{z^{N_{1}+\deg(a_{n,j})}}\right)\right)f=0,$ (4.5)
@@
$P_{i,j}(z)=z^{d_{i,j}}\sum_{v=0}^{\infty}\frac{b_{i,j,v}}{z^{v}},\quad|z|>|z_{0}|,$ (4.6)
for $d_{i,j}\in\mathbb{N},b_{i,j,v}\in\mathbb{C},i=1,\ldots,n-1,$ $j=0,\ldots,N_{0}^{\prime},$ enough large $|z_{0}|$, and $z\in S_{j}(\varepsilon)$. $Q_{0}$ is consequently in the asymptotic form of
$Q_{0,j}(z)\sim z^{d_{0,j}}\sum_{v=0}^{\infty}\frac{b_{0,j,v}}{z^{v}},\quad|z|>|z_{0}|,$ (4.7)
@@
Next, we introduce the essence of $|z|\in E$. In the proof of Gundersen’s result [13, Corollary 3], the essential idea is to use Cartan’s Lemma [13, Lemma 9] to dig out constructive disks containing the zeros and poles of $f^{(j)}$ (He defaults $f$ as a meromorphic function and here we consider $j=0$ and $f$ entire). And introducing the set of exceptions $E$ with respect to $r$ is exactly the sum of the diameters of these disks. We recall the constructive part of this [13, pp. 100-101], which begins by partitioning the complex plane into a merge of countably many annular strips, i.e., there exists an integer $n_{0}>0$, and for an arbitrary integer $\nu\geq n_{0}$ and a constant $\alpha>1$, there is
$\alpha^{\nu}\leq|z|\leq\alpha^{\nu+1}.$ (4.8)
In each such strip, there will be only a finite number of zeros of $f$ and closed disks $B_{\nu,1},\ldots,B_{\nu,p_{\nu}}$ containing them (note that the center of the circle here is not necessarily a zero of $f$, and a single disk may contain more than one zero). For any $\varepsilon_{0}>0$, by [13, p. 101], when we take the sum of the diameters of these disks not exceeding $4\alpha^{-\varepsilon_{0}\nu}$ the corresponding $E$ is obtained as a finite linear measure by Cartan’s Lemma [13, Lemma 9].
Thus, (4.5) can be modified as
$f^{(n)}+P_{n-1,j}f^{(n-1)}+\cdots+P_{1,j}f^{\prime}+\left(P_{0,j}+o\left(\frac{1}{z^{N_{1}+\deg\{a_{n,j}\}}}\right)\right)f=0,$ (4.9)
which is established for $z\in S_{j}^{*}$, i.e., $z\in S_{j}(\varepsilon)$ and
$z\notin\tilde{B}:=\bigcup_{\nu=n_{0}}^{\infty}\bigcup_{t=1}^{p_{\nu}}B_{\nu,t},~n_{0}>N.$
In a step further, it can be shown that (4.9) satisfies the form of (2.9) if $f$ only possesses finite number of zeros in $S_{j}(\varepsilon)$, which means $\tilde{B}$ disappears when $|z|>|z_{0}|$ for sufficiently large $|z_{0}|$. Using Lemma 2.4, there exists a polynomial $G_{\theta}$ in $z^{1/p_{j}}$, $p_{j}\in\mathbb{N}$, and a set of zero relative measure $E_{\theta}$ such that